OpenStudy (anonymous):

Linda runs 7 miles in 60 minutes. At the same rate,how many miles would she run in 24 minutes?

OpenStudy (ajprincess):

for 60 minutes she runs 7 miles for 24 minutes say she runs x miles. 60:7=24:x 60/7=24/x solve for x:)

OpenStudy (anonymous):

i got 2.5 as the answer

OpenStudy (ajprincess):

nope:(

OpenStudy (anonymous):

well i don't understand then :l i took 60/24 and got 2.5, i don't know how to solve for x D:

OpenStudy (ajprincess):

2.5=7/x 2.5/1=7/x cross multiply x=7/2.5

OpenStudy (anonymous):

wait, i have to cross multiply right? like proportions

OpenStudy (ajprincess):

ya:)

OpenStudy (anonymous):

is the answer gonna be 2.8m?

OpenStudy (ajprincess):

ya:)

OpenStudy (anonymous):

thanks!